OpenStudy (anonymous):

Which equation is an example of a direct variation? A) y=2/3xz B) 2xy=12 C) y=8/x D) 2/3x=y

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hint: a direct variation is y = kx where k is any real number

OpenStudy (anonymous):

c?

jimthompson5910 (jim_thompson5910):

that's an inverse variation equation

jimthompson5910 (jim_thompson5910):

since it's in the form y = k/x

OpenStudy (anonymous):

ohhhh so a?

jimthompson5910 (jim_thompson5910):

that's a joint variation equation, notice you have 2 variables on one side

OpenStudy (anonymous):

hmm b?

jimthompson5910 (jim_thompson5910):

no lol, sounds like you're just guessing at this point?

jimthompson5910 (jim_thompson5910):

choice B is another inverse variation equation, just in a different form

OpenStudy (anonymous):

ok d lol . and i thought d wasnt it cause of its form .

jimthompson5910 (jim_thompson5910):

2/3x=y y = 2/3x its in the form y = kx where k = 2/3

OpenStudy (anonymous):

oh ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome